how would i do this question
? Chris has a basket of apples, 40% of which are red. The rest of the apples are green. If the basket contains 6 red apples, how many green apples does it contain?

Answer by nyc_function(2741)

0.40/0.60 = 6/x


0.60 times 6 = 0.40x


3.6 = 0.40x


3.6/0.40 = x


9 = x


Nine are green.